OpenStudy (anonymous):

the length of a rectangular garden is 4m more than its width. of the area of the garden is 124 m^2, find the length and width of the garden

OpenStudy (misty1212):



OpenStudy (misty1212):

\[x(x+4)=124\\ x^2+4x=124\\ (x+2)^2=128\]